OpenStudy (anonymous):

so what is the sum of 1/3 and 1/6

OpenStudy (karatechopper):

ok well what you do, is find a common denominator...so can u think of one that bothh 3 and 6 can go into?

OpenStudy (lgbasallote):

1/3 + 1/6 = (6+3)/(18) = 9/18 = 1/2

OpenStudy (unklerhaukus):